### v1

# OSPTrack: A Labeled Dataset Targeting Simulated Open-Source Package Execution

Abstract
Open-source software is a fundamental part of the internet and the cyber supply chain, but its exploitation has become more frequent. While vulnerability detection in OSS has advanced, previous work mainly focuses on static code analysis, neglecting runtime indicators. To address this, we created a dataset spanning multiple ecosystems, capturing features generated during the execution of packages and libraries in isolated environments. The dataset includes 9,461 package reports (1,962 malicious), with static and dynamic features such as files, sockets, commands, and DNS records. Labeled with verified information and detailed sub-labels for attack types, this dataset helps identify malicious indicators, especially when source code access is limited, and supports efficient detection methods during runtime.
Index Terms:
Open-source Software, Sandbox Simulation, Static Features, Dynamic Features, Supply Chain Security
I Introduction
Open source software (OSS) has become a core component of a wide range of applications, but its integration has introduced vulnerabilities that are increasingly exploited in adversary activities [1]. Existing vulnerability detection methods often focus on static code analysis [2, 3], which can be inefficient when OSS is embedded within complex systems. Current datasets typically focus on malicious packages [4, 5, 6], but fail to capture runtime features. Although software simulations have been developed, no dataset specifically targets OSS and provides labeled monitoring results during execution.
This dataset leverages the package-analysis tool  [^1] https://github.com/ossf/package-analysis to simulate package execution, using authenticated references from existing datasets in the supply chain security domain [7, 5]. By parsing the generated reports and extracting features, we provide a comprehensive labeled dataset with eight-dimensional features, including files, sockets, commands, and DNS-related behaviors.
The dataset, generated over five weeks, includes 9,461 valid items across ecosystems such as npm, pypi, crates.io  [^2] https://crates.io/, nuget  [^3] https://www.nuget.org/, and packagist  [^4] https://packagist.org/. The data, along with the code for simulating package execution, feature extraction, and label matching, is available online [8] to facilitate reproducibility and further research.
II Background
II-A OSS Supply Chain Security
Ladisa et al. [2] identified the means of spreading malware through the Open Source Software (OSS) supply chain. By analyzing Backstabber’s Knife Collection [9], they uncovered several exploitation cases, including reverse shell spawning a shell process, dropper with second-stage payload, and data exfiltration, among others.
Zhang et al. [10] present a categorization of detection methods for malicious packages. They identify three main categories: rule-based, unsupervised, and supervised learning approaches. Rule-based methods prioritize precise matching of IOCs associated with malicious behaviors, such as specific IP addresses, domain names, or registry values. This method is widely used in industry because of its accuracy.
Liang et al. [3] explored the detection for mobile applications. The proposed method created graph representation of code property graphs (CPGs) from source code and the detection combines advanced machine learning algorithms like CodeBERT, and attention of Transformers with Graph Convolutional Networks (GCN).
Tran et al. [6] introduced a BERT-based architecture to detect vulnerable statements in Python code. This framework is capable of learning complex relationships between code statements without predefined graph structures.
The majority of current work on OSS vulnerabilities [11, 12, 13] focuses on plain code and functions, or dependencies. For running-status behaviors, the related works focus more on malware or ransomware detection [14, 15], rather than malicious scripts in OSS.
II-B Threat Data Generation
Zhang et al. [10] described that learning-based approaches focus on extracting discrete features, which can be obtained through simulation, dynamic, or static analysis. Furthermore, three execution scenarios have been highlighted for triggering malicious behavior: install-time, import-time, and run-time execution. These scenarios inspire us to perform appropriate simulation methods to encompass these behaviors.
Landauer et al. [16] performed full cyber kill chain in their attack simulation. They carried out reconnaissance via port and vulnerability scanning, delivery through uploaded webshell for exploitation, and etc, to obtain root privilege for command execution. The simulation of full cyber kill chain released more comprehensive exploitation behavior. However, due to the design complexity, it is not possible to contain too many scenarios, which reduced the potential generality.
Parisot et al. [14] introduced a novel approach for ransomware detection by combining the dynamic analysis features of the Cuckoo Sandbox  [^5] https://cuckoosandbox.org/ with machine learning techniques. Their method involved creating three datasets that capture API calls, network interactions, and string data from malware samples, with feature extraction using the TF-IDF technique [17].
Silva et al. [15] employed dynamic analysis and machine learning to identify ransomware signatures using selected dynamic features. The research involves experiments with encryptor and locker ransomware, alongside goodware, to generate a dataset of dynamic parameters in a controlled environment using a sandbox.
Regarding OSS, some datasets have been generated and released for research and defense purposes [4, 5, 7]. However, these datasets primarily focus on collecting malicious packages and libraries, rather than analyzing their runtime behaviors. While a public dataset has been generated using the tool package-analysis, it lacks detailed labeling and specific feature information, which significantly limits its utility.
III Methodology
To generate a labeled OSS-tailored dataset focusing on running-time features, we adopt the methodology demonstrated in Fig. 1. The core design includes multi-process analysis for packages, report parsing with feature extraction, and final label matching. Additionally, part of dataset, espesically majority of benign examples, is from the queried results from available BigQuery dataset generated by package-analysis tool. The detailed pipelines are illustrated below.
Figure 1: The Data Generation Framework. 1. Collect package information 1.a. Query analyzed results from BigQuery 2. Simulate packages with package-analysis in multiple processes 3. Parse JSON reports and also queried Parquet reports, extract features 4. Match and generate labels based on known labels
III-A Package Collection
When collecting malicious package information, two datasets are included, which are Backstabbers-Knife-Collection (BKC)[9] and malicious packages repository [7]. The first dataset with a total number of 5942, covering npm, pypi, and ruby, has been widely referred in academia [11, 2]. The second one covers the majority of revealed supply chain attacks until November 2024. This dataset is made by the open source security foundation, which is the most comprehensive dataset in the OSS ecosystem. This dataset has a total number of 24215 samples and contains more ecosystems with crates.io, nuget, and packagist.
The malicious packages in two datasets are organized with JSON files to introduce package information and potential source code if available. Inside per JSON file, the included information is package name, version, ecosystem, indicators of compromise (iocs), and details regarding exploitation method. Based on the core indicators (blockchain, source code, commands, etc) in description, we further define the potential attack type, malicious functions, trigger mechanism, code location, and evasion method referring to [9, 10].
After aggregating all information, two corresponding files are generated, which are named data_iocs.csv and data_metrics.csv. The iocs file collects all the indicators spanning domains, IP addresses, links, urls, and strings, which can be taken as the set to directly identify malicious packages or libraries. The metrics file integrates the details of malicious packages, which can be used for exploring the high-level package behaviors and demonstrating the tendency for package exploitation in diverse ecosystems. The metrics file has been further used for label and sub-label matching, due to the specification of package information. The statistical information from the metrics file is provided in Table I.
TABLE I: Collection of Attack Types in Software Supply Chain
Attack Type Number
data exfiltration/root shell 12026
dropper/malware 5425
typosquatting 937
C2 communication 1674
malicious command execution 118
starjacking 31
social engineering/typosquatting 14
typosquatting/malware 1
A portion of the dataset comes from the BigQuery dataset generated by the package-analysis tool. This public dataset is based on live instances of newly published packages or libraries, with reports that include both static and dynamic features. However, the dataset does not include explicit labels for the reports, so labeling still needs to be confirmed.
In total, we collect 7,500 package reports from this dataset, which constitutes the majority of the benign samples. These reports are saved in Parquet  [^6] https://parquet.apache.org/ format to preserve the original data types, such as lists or dictionaries, rather than converting them to strings as in CSV format. This approach eliminates the need for additional transformations and complex string parsing operations.
III-B Simulation
Package-analysis is a type of sandbox as the core that can simulate the package execution. Sandbox has been widely adopted for analyzing dynamic behaviors of software and detecting whether a software is malicious [18, 19]. Some known sandbox examples include Cuckoo and VirusTotal [20]. Different from those sandboxes, package-analysis is capable of inputting package info like ecosystem, name, and version, output of real code or packages for analysis.
The analysis results generated by this tool are the extracted dynamic and static analysis implemented through simulated execution of malicious package. The dynamic analysis inside those results is complete HTTP or HTTPS dynamic analysis collection. To further explain, it combines the information of DNS query, command execution, socket connection, standard errors (stderr), file manipulations, status, and standard output (stdout). In contrast, the static analysis collects the direct information like suspicious identifiers, escaped strings, based64 strings, hex strings, IP addresses, URLs, and etc.
The scheduler in this tool creates analysis jobs from Package Feeds. The analysis component performs one-time tasks, with workers collecting package behavior data through both static and dynamic analysis. By default, results are pushed to BigQuery for public access. To store results locally, we modified the simulation script to save files in the format ecosystem-package-version. To prevent duplicates, we also implemented a check for previously generated results.
The simulation runtime varies from several seconds to over 20 minutes, depending on the size of the analyzed package. Some packages, as recorded in [7] and [5], failed to be analyzed due to missing dependencies. Furthermore, certain packages caused the simulation to become stuck, affecting the analysis of subsequent packages in the queue. To address this, we implemented a timeout mechanism (e.g., 10 minutes) and regular health checks to detect and mitigate simulation stalls.
In terms of performance, a single process generates an average of 50 package reports per day. By implementing multiple processes, each starting from different row indices, the generation speed increases to approximately 150 packages per day.
In five weeks, 4235 reports have been created, in which only 1961 is valid. We can verify the large proportion of insufficient simulations. The main reason behind this is missing, mitigated, or removed historical packages after identifying vulnerabilities. Those reports are in JSON format and processed in the next stage for parsing and feature extraction.
III-C Feature Extraction
After collecting both reports from package-analysis and BigQuery dataset, the next pipeline is the parsing engine, in which feature extraction is included. The feature extraction involves extracting package information with package name, version, and its belonging ecosystem from Package key, and the rest detailed features from system and network interactions, which include execution, import, and install behaviours. Under both import and install sections, the collected features include Status, Stdout, Stderr, File, Sockets, Commands, and DNS. However, the values of some keys are encoded random code (Stdout, Stderr) or contain non-important information (Status). After bypassing those unnecessary features, we choose only remaining four-dimensional features, distributing under two categories (import, install), resulting in eight features out of package-related features.
III-D Labeling Method
The labeling process has been divided into two separate parts, in order to process separate data from simulation and BigQuery. As demonstrated in Algorithm 1, The labeling has also included two layers, in which the first layer defines malicious(1) or benign(0) and the second layer denotes specific attack types under malicious category. The matching references are generated through two available dataset [9, 7], denoted as bkc_mal.csv and pkg_mal.csv in dataset.
Algorithm 1 Label Matching Algorithm
$D_{1}$, $D_{2}$, $R_{f}$ (Matching Reference), $M_{f}$ (Metric)
2:Combined Dataset with Updated $L$ (Labels) and $SL$ (Sub-Labels)
3:Initialize: Load $D_{1}$, $D_{2}$, $R_{f}$, and $M_{f}$.
4:for all records in $D_{1}$ do
5:     Extract $(P,E,V)$ from record. ($P$: Name, $E$: Ecosystem, $V$: Version)
6:     Search $R_{f}$ for matching $(P,E,V)$.
7:     if match found then
8:         Update $L$ in $D_{1}$ using $R_{f}$.
9:     end if
10:end for
11:for all records in $D_{2}$ (default $L=1$do
12:     Extract $(P,E,V)$ from record.
13:     Search $M_{f}$ for matching $(P,E,V)$.
14:     if match found then
15:         Assign $SL$ using attack_type in $M_{f}$.
16:     end if
17:end for
18:Combine $D_{1}$ and $D_{2}$, preserving all records and updated $L$, $SL$.
19:Output: Combined dataset with updated $L$ and $SL$.
IV Dataset Description
The entire simulation process and experiments are conducted on a virtual machine running Ubuntu 22.04.5 LTS with 8 gigabytes of memory. The generated labeled dataset includes 7499 benign examples and 1962 malicious examples, spanning five ecosystems (pypi, ruby, npm, rust, php). The dataset is stored in pickle  [^7] https://docs.python.org/3/library/pickle.html format to remain the original data type. The distribution of labels has been illustrated in Table II, from which we can see the majority of packages is still on npm ecosystem and ruby takes least proportion.
TABLE II: Package counts by ecosystem, package count, label, and sub-label.
Ecosystem Count Label Sub_Label
crates.io 1205 0 na
npm 4645 0 na
packagist 265 0 na
pypi 1323 0 na
rubygems 61 0 na
crates.io 1 1 na
npm 800 1 na
pypi 812 1 na
rubygems 269 1 na
npm 18 1 C2
pypi 38 1 C2
rubygems 8 1 C2
npm 11 1 root shell
npm 2 1 command exec
pypi 2 1 command exec
Apart from the main labeled dataset, we also include extracted files from available datasets. Those files can be further classified into statistical data and information data. The first group of statistical data denote the iocs file and metrics file created from dataset [7]. They are used for matching sub labels and statistic analysis on malicious behaviors. The information data denotes the package information collected from two datasets, playing as references to match malicious labels for queried data.
V Research Opportunities
The dataset is primarily designed for supply chain security research. It can be used to train machine learning models to differentiate between benign and malicious software behaviors, as well as detect vulnerabilities in running software to ensure supply chain security in open-source software (OSS).
Additionally, it is possible to identify indicators triggered by malicious activity, which can aid in detection and prevent further exploitation. The dataset’s diversity also enables differential or comparative analysis of malicious package behaviors across different ecosystems [21].
Moreover, more detailed feature extraction can be performed on this dataset, allowing for better attack behavior profiling [22] and facilitating representation learning, such as graph-based representations [23, 24].
VI Limitations
Since the source code is unavailable, the simulated scenarios cannot fully capture the injection process at the source due to the wide variety of exploitation methods. Additionally, some malicious packages could not be analyzed or have reports generated due to missing or deleted packages. Packages requiring extended analysis time were excluded from the dataset because of timeout settings. Furthermore, only a portion of the available malicious packages was simulated due to time constraints. We plan to regularly update the dataset to include a more extensive and diverse collection of malicious reports.
VII Conclusion
This work presents a running-time-oriented labeled dataset for supply chain security research, with ground-truth labeling and sub-labels for attack types. It covers ecosystems such as npm, Packagist, PyPI, RubyGems, and crates.io, and includes eight extracted features representing static and dynamic behaviors. The dataset is suitable for vulnerability detection and malicious software classification. Future work will focus on expanding the dataset with additional examples for specific ecosystems and analyzing its application for vulnerability detection.
Acknowledgments
This research has received funding from JUMPSEC Ltd.